OpenStudy (anonymous):

solve sqrt 9x+55=x+5

OpenStudy (ash2326):

flycherry u should try this by yourself, if ur stuck somewhere i'll help u

OpenStudy (ash2326):

did u try flycherry??

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

no i dint try because i dont know where to start lol

OpenStudy (ash2326):

ohk we have\[\sqrt{9x+55}=x+5\] now square both sides

OpenStudy (anonymous):

so I am squaring the 9 and the 5 at the end?

OpenStudy (anonymous):

thats 81 and 25

OpenStudy (ash2326):

no. square both sides

OpenStudy (anonymous):

i dont know what that means

OpenStudy (ash2326):

left hand side has a root so if u square it u are left with 9x+55

OpenStudy (anonymous):

that makes absolutely no sense tome

OpenStudy (ash2326):

now on the right hand side we have (x+5) you know (a+b)^2=a^2+b^2+2ab apply this to right hand side

OpenStudy (anonymous):

if you are squaring a root then it would be 2sqrt then the rest... lol

OpenStudy (ash2326):

nope

OpenStudy (ash2326):

square root of 4 is 2 if you square 2 you get 4 again , did u understand

OpenStudy (ash2326):

square of a square root term and we r left with only the term

OpenStudy (ash2326):

are u getting my point

OpenStudy (anonymous):

omg this same question is on my test

OpenStudy (ash2326):

see our question is \[\sqrt{9x+55}=x+5\] now square right hand side as well as the left hand side we get \[9x+55=(x+5)^{2}\]

OpenStudy (ash2326):

now use the formula for (a+b)^2=a^2+b^2+2ab

OpenStudy (ash2326):

so we get \[9x+55=x^2+5^2+10x\]

OpenStudy (ash2326):

\[9x+55=x^2+25+10x\] bring all the terms to left hand side we get 0=x^2+x-30

OpenStudy (ash2326):

now factor the term 30,such that both factors differ by 1 5,6 are the two factors of 30 satisfying this

OpenStudy (ash2326):

so x can be written as 6x-5x

OpenStudy (ash2326):

our equation can be rewritten as x^2+6x-5x-30=0

OpenStudy (ash2326):

now take out the common factors x(x+6)-5(x+6)=0 takin out (x+6) common we get (x+6)(x-5)=0 so x=-6 or x=5

OpenStudy (ash2326):

flycherry did u understand?